A particle carrying +2.35 \({\rm \; nC}\) of positive charge is at the origin of an xy coordinate system. Take \(V(\infty )\) to be zero for these calculations.
Calculate the potential \(V\) on the \(x\) axis at \(x\) = 3.0000 \(\rm m\).
1
Expert's answer
2019-09-16T09:45:54-0400

Potential, created by a point charge "q" at a distance "r" from the charge is "V_E = \frac{1}{4 \pi \varepsilon_0} \frac{q}{r}" , where "\varepsilon_0 = 8.85 \times 10^{-12} F \cdot m^{-1}" is the vacuum permittivity.

Hence, "V_E= 9 \cdot 10^9 N \cdot m^2 \cdot C^{-2} \times \frac{2.35 \cdot 10^{-9} C}{3 m} = 7.05 V".